what mass of glucose C6H12O6 must be dissolved in 455g ethanol C2H5OH to prepare a 2.40x10^(-2)

m solution?


1
Expert's answer
2021-04-21T03:53:20-0400

Molarity = moles /solution

= 2.40 = Y/X + 563

= 2.40(X+563)

=2.40 X + 13151.2

Molar Mass of glucose =180.156

= 180.156/13151.2

= 2.43g